OpenStudy (anonymous):

Find the solution of each of the following differential equations for a function x = x(t): dx/dt − 2(x/t) = t^2 − 1 (Hint: Find an integrating factor) Your assistance is appreciated. Kind regards

OpenStudy (anonymous):

My work so far: 1) integrating factor , using -2/t is : e^integral(-2/t dt 2) e^integral(-2/t dt = e^-2lnt = -2t hence integrating factor is -2/t , or is it -2/t Correct so far?

OpenStudy (anonymous):

\[ \Large \frac{dy}{dt}+ay=g(t) \\ \Large \mu(t)=\exp\left(\int adt\right) \]

OpenStudy (anonymous):

Pardon me, I am writing this down for myself basically to remember things, but you agree with this so far?

OpenStudy (anonymous):

just one second,,, yes

OpenStudy (turingtest):

yup, that's correct

OpenStudy (anonymous):

So if you do some pattern matching it seems to me like you have to do the following \[ \Large \mu(t) =\exp \left( -\int 2tdt\right)\]

OpenStudy (anonymous):

is that 2t or 2/t

OpenStudy (anonymous):

It is 2t here, that's how I would have done it anyway.

OpenStudy (anonymous):

agree @TuringTest ?

OpenStudy (anonymous):

Sorry , the question was messed up when I copied and pasted it. That shoul be -2(x/t)

OpenStudy (turingtest):

-2t yes'

OpenStudy (turingtest):

oh, well...

OpenStudy (anonymous):

ohh well that changes things hehe.

OpenStudy (anonymous):

I will edit the question...my apologies

OpenStudy (turingtest):

then let's rewrite the question it's okay, it happens

OpenStudy (anonymous):

hehe it certainly does hehe

OpenStudy (anonymous):

fixed!

OpenStudy (anonymous):

\[\Large \frac{dx}{dt}- \left( \frac{2}{t}\right)x=t^2-1 \]

OpenStudy (anonymous):

so your integration factor is correct

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

which one is correct is that 2t or 2/t

OpenStudy (turingtest):

as you said 2) e^integral(-2/t dt = e^-2lnt = -2t that's your IF

OpenStudy (anonymous):

\[ \Large \exp\left(-2\int \frac{1}{t}dt=-2\ln(t) \right) \\ \Large e^{-\ln t 2}=-2t\]

OpenStudy (turingtest):

oh your last par was wrong

OpenStudy (anonymous):

okay....I was getting -2/t from wolframalpha , so i got confused, because my own work gives me 2t...

OpenStudy (turingtest):

\[e^{-2\ln t}=e^{\ln(t^{_-2})}=t^{-2}\]

OpenStudy (anonymous):

yes I just noticed that I was making a bad substitution there, my bad.

OpenStudy (turingtest):

funny how you guys both made the same mistake though :p

OpenStudy (anonymous):

I should have written it out with exponentials (-:

OpenStudy (anonymous):

lol so the correct result is t^-2

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

..for the integrating factor

OpenStudy (anonymous):

thanks....I will attempt the solution by myself in a bit ...If I need any help I will give you gentlemen a shout...:)

OpenStudy (anonymous):

thank you so much

OpenStudy (turingtest):

welcome!